OpenStudy (anonymous):

The equation f(x) = 3x2 − 24x + 8 represents a parabola. What is the vertex of the parabola? (−4, 152) (4, −40) (−3, 107) (3, −37)

OpenStudy (anonymous):

i dont know what a parabola is

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

a symmetrical open plane curve formed by the intersection of a cone with a plane parallel to its side. The path of a projectile under the influence of gravity ideally follows a curve of this shape.

OpenStudy (danjs):

you could complete the square to make the vertex equation form

OpenStudy (anonymous):

ok i know how to complete the sqaure

OpenStudy (anonymous):

8, 1\3

OpenStudy (danjs):

y = a(x-h)^2 + k vertex at (h,k)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

ive never heard of that

OpenStudy (anonymous):

this is so confusing

OpenStudy (danjs):

the video that was linked is good

OpenStudy (anonymous):

ok ill watch it real quick brb

OpenStudy (danjs):

If you havent done anything else yet , just plug in the numbers to that

OpenStudy (anonymous):

yea i dont get it

OpenStudy (danjs):

\[y = ax^2 + bx +c\] x coordinate of the vertex will be \[\frac{ -b }{ 2a }\]

OpenStudy (danjs):

b = -24 a = 3

OpenStudy (anonymous):

so divide that?

OpenStudy (danjs):

yes, just calculate it, put in the , a and b, values

OpenStudy (anonymous):

ok ill do that now

OpenStudy (anonymous):

i got -8b\a

OpenStudy (danjs):

X - Coordinate of Vertex \[\frac{ -b }{ 2a } = \frac{ -(-24) }{ 2*3 } = \frac{ 24 }{ 6 } = 4\]

OpenStudy (anonymous):

i have to leave in like 3 minutes should i just guess??

OpenStudy (anonymous):

wait its b ???

OpenStudy (danjs):

The X coordinate for the vertex is -b/(2a) = 4 To get the Y-Coordinate for that X value, plug in x=4 into the original y = 3x^2 - 24x + 8

OpenStudy (danjs):

well you dont have to do that, if the X coordinate is 4, there is only one choice

OpenStudy (anonymous):

i just turned it in, b is right :)

OpenStudy (danjs):

yea just have to remember the \[\frac{ -b }{ 2a }\] for the x-coordinate of the vertex